Jefferson Lab HALL A Logbook for December 07 - All EXCEPT Automatic Entries

  • Main INDEX

    Monday

  • 222447 12/31/07 21:31 A. Camsonne Pedestal channel 254 reduced
  • 222445 12/31/07 16:18 jpchen right VDC bottom started flow
  • 222444 12/31/07 16:09 R. Michaels EDTM delay offsets
  • 222443 12/31/07 16:08 jpchen Detector gas system alarm.
  • 222439 12/31/07 15:20 atrig Trigger-Download
  • 222435 12/31/07 11:21 A. Camsonne Gas alarm counting house
  • 222434 12/31/07 11:08 A. Camsonne Right arm iocha16 rebooted
  • 222433 12/31/07 10:25 Ramesh No more fire alarm
  • 222432 12/31/07 10:04 Ramesh Fire alarm
  • 222430 12/31/07 00:35 A. Camsonne Left arm pedestal subtraction back on
  • 222427 12/31/07 00:20 A. Camsonne NaI HV on for cosmics
  • Saturday

  • 222426 12/22/07 11:26 Ramesh Four cables from the Hall patch pannel unplugged
  • Thursday

  • 222424 12/20/07 13:30 folts diffuser/tech
  • 222423 12/20/07 11:32 folts holiday shutdown
  • 222422 12/20/07 10:47 A. Camsonne Pedestal subtraction off for left arm
  • 222410 12/20/07 09:45 A. Camsonne NaI HV off
  • 222409 12/20/07 08:37 Ramesh HRS magnets ramped down, detector's HV turned off
  • 222408 12/20/07 08:17 V. Sulkosky Boiling Test Run Summary
  • 222406 12/20/07 08:07 V. Sulkosky Owl Shift Summary
  • 222405 12/20/07 08:05 gilman target charts just after beam ended
  • 222384 12/20/07 07:24 V. Sulkosky Radiation/Ozone chart
  • 222373 12/20/07 06:25 V. Sulkosky Raster/BPM Plots
  • 222364 12/20/07 05:21 V. Sulkosky Replay Macros Modified to Write to adaql10
  • 222363 12/20/07 05:15 V. Sulkosky BCM and BPM Strip Charts
  • 222358 12/20/07 04:54 V. Sulkosky Out of Disk Space for Replay
  • 222355 12/20/07 04:39 V. Sulkosky Compton/OTR Charts
  • 222350 12/20/07 04:14 V. Sulkosky LHRS and RHRS Collimator OPEN
  • 222345 12/20/07 04:01 V. Sulkosky Beam Position and Raster
  • 222344 12/20/07 04:01 Holmstrom replay failing
  • 222343 12/20/07 03:57 V. Sulkosky RHRS sieve slit IN
  • 222338 12/20/07 03:15 V. Sulkosky LHRS Sieve Slit IN
  • 222332 12/20/07 02:12 Holmstrom Maybe Just a lot of Pions
  • 222331 12/20/07 02:05 Holmstrom Shower Calibration?
  • 222322 12/20/07 01:33 V. Sulkosky RHRS Dipole Search Off
  • 222314 12/20/07 00:21 techs IOC
  • 222313 12/20/07 00:20 V. Sulkosky Hall A Tools
  • 222312 12/20/07 00:14 T. Ewing/A. Left Arm slow controls problems solved
  • 222311 12/20/07 00:04 MinaKat Shift Summary
  • Wednesday

  • 222309 12/19/07 23:49 minakat Hall A Tools screen
  • 222308 12/19/07 23:44 Makis P TGT Controls/Positions
  • 222307 12/19/07 23:42 Makis P Loops charts
  • 222306 12/19/07 23:40 Makis P optics 7-foil tgt in
  • 222305 12/19/07 23:33 minakat Left IOC/run plan update
  • 222304 12/19/07 22:44 A. Camsonne Right arm NMR lock
  • 222303 12/19/07 21:45 A. Camsonne Lost spectrometer IOCs
  • 222298 12/19/07 21:13 A. Camsonne Target moved to Loop3 H2
  • 222295 12/19/07 21:05 A. Camsonne NaI HV change
  • 222292 12/19/07 20:45 A. Camsonne Target moved from Loop 3 H to Optics
  • 222290 12/19/07 20:39 A. Camsonne NaI HV change
  • 222287 12/19/07 20:31 Makis P H2-15cm-Lp3 IN, switched coolant
  • 222263 12/19/07 19:23 minakat spot++
  • 222248 12/19/07 19:03 Makis P He10cmLoop1 in Beam
  • 222235 12/19/07 18:30 jpchen Reduced loop 2 coolant and fan speed
  • 222214 12/19/07 17:56 Makis P C12 tgt IN
  • 222213 12/19/07 17:48 A. Camsonne NaI HV change
  • 222200 12/19/07 17:23 A. Camsonne NaI HV change
  • 222179 12/19/07 16:41 minakat spot++
  • 222174 12/19/07 16:29 Makis P He10cmLoop1 in Beam
  • 222173 12/19/07 16:22 Ramesh BCM calibration complete
  • 222169 12/19/07 16:00 X. Zheng day shift summary
  • 222166 12/19/07 15:24 Ramesh Hall A tools ready for BCM calibration
  • 222165 12/19/07 14:56 Jack Ethane cylinder seems ok
  • 222164 12/19/07 14:19 Aniol r4308 no histos in replay
  • 222163 12/19/07 13:58 Aniol target moved to empty
  • 222144 12/19/07 12:26 Aniol reading target pressures
  • 222141 12/19/07 12:22 X. Zheng BCM screen
  • 222140 12/19/07 12:21 X. Zheng Ozone chart
  • 222139 12/19/07 12:20 X. Zheng Hall A Tools page
  • 222126 12/19/07 11:23 fansler Left Q1 Trips on He level
  • 222125 12/19/07 11:19 Aniol r4299 replay has no histograms
  • 222124 12/19/07 11:17 Jack Left Q1 reset
  • 222123 12/19/07 10:59 aniol 4301 has no hist, root file split?
  • 222122 12/19/07 10:32 Aniol split root files, 4296,4298
  • 222121 12/19/07 10:22 X. Zheng controlled access to reset Left Q1
  • 222117 12/19/07 10:01 X. Zheng Left Q1 trip
  • 222107 12/19/07 09:38 X. Zheng Right arm start-run comment wrong for collimator since run 21982
  • 222100 12/19/07 09:20 Aniol root files 4292, 4293 split
  • 222096 12/19/07 08:54 Aniol target charts
  • 222086 12/19/07 08:08 johna Owl shift summary/runlist
  • 222064 12/19/07 06:51 Kiadtisak TS0 failed
  • 222058 12/19/07 06:25 johna Changed HRS-L momentum
  • 222047 12/19/07 05:44 johna Snapshot: BCM/BPM screens
  • 222046 12/19/07 05:43 johna Snapshot: Tools and ozone/rad monitor screens
  • 222030 12/19/07 04:32 johna Changing spectrometer momenta
  • 221996 12/19/07 02:34 johna Changing HRS-L momentum
  • 221978 12/19/07 01:30 Yoomin Target chart
  • 221971 12/19/07 00:45 johna Changing HRS momenta
  • 221957 12/19/07 00:02 Makis P Shift Summary
  • Tuesday

  • 221955 12/18/07 23:58 minakat target is Pb loop 2
  • 221950 12/18/07 23:49 minakat target is He loop 2
  • 221949 12/18/07 23:43 Makis P Ozone/rad Chart
  • 221948 12/18/07 23:41 Makis P Current Monitor
  • 221947 12/18/07 23:40 Makis P Hall A Tools
  • 221942 12/18/07 23:27 minakat target is Fe Left
  • 221939 12/18/07 23:24 Makis P Bream energy reading OK now
  • 221936 12/18/07 23:12 minakat target moved from dummy 10cm to C12 left
  • 221933 12/18/07 23:01 Makis P Cannot run at 20uA
  • 221928 12/18/07 22:49 Makis P Offline Note/Beam Energy
  • 221925 12/18/07 22:44 minakat target set: H2 loop 2
  • 221922 12/18/07 22:32 Makis P Beam x-position set at -3.5 BPMA/B
  • 221919 12/18/07 22:24 Makis P Finished Pb tgt beam x-scan
  • 221918 12/18/07 22:06 A. Camsonne Pb position scan
  • 221917 12/18/07 22:03 A. Camsonne Pb target position scan
  • 221872 12/18/07 20:01 minakat target in : Pb Loop3
  • 221869 12/18/07 19:53 Makis P Finished unrastered beam scan/Empty tgt
  • 221866 12/18/07 18:43 minakat target is empty
  • 221859 12/18/07 18:30 Jack Measurements on Q2's and Q3's during access
  • 221858 12/18/07 18:27 Jack (for He Left Q1 He level readback
  • 221855 12/18/07 18:25 jpchen Run 4232 has target x between -1 to -2 mm.
  • 221826 12/18/07 17:41 minakat target set : Pb loop 3
  • 221825 12/18/07 17:40 Makis P Spot++ Empty tgt/Beam looks fine
  • 221822 12/18/07 17:32 jpchen Target: changed loop 12 to 7K, 48 psi.
  • 221821 12/18/07 17:32 Makis P Beam back, 5uA, do Pb raster scans
  • 221816 12/18/07 17:24 Makis P Harp Scans
  • 221813 12/18/07 16:25 minakat target charts-no beam
  • 221812 12/18/07 16:02 Holmstrom Day Shift Log
  • 221810 12/18/07 14:34 Aniol hall controlled access and LQ1 trip
  • 221809 12/18/07 12:34 Aniol run 4232 lead scan -0.5mm
  • 221808 12/18/07 12:26 Holmstrom momentum change
  • 221807 12/18/07 09:25 Aniol lead scan +2mm, +3mm
  • 221806 12/18/07 09:22 Aniol lead scan 0mm, +1mm
  • 221805 12/18/07 09:19 Aniol lead scan, -2mm and -1mm
  • 221804 12/18/07 08:58 Aniol empty target in position
  • 221803 12/18/07 08:57 Aniol run 4031,+4mm Pb scan
  • 221802 12/18/07 08:43 Holmstrom Beam Away
  • 221800 12/18/07 08:30 A. Camsonne Moving replayed files to adaql10
  • 221792 12/18/07 08:14 Yoomin Target alarm
  • 221790 12/18/07 08:06 johna owl shift summary/list of runs
  • 221784 12/18/07 07:46 johna ozone and radiation monitor screen
  • 221782 12/18/07 07:45 johna Tools screen and BCM/BPM
  • 221771 12/18/07 06:42 R. Michaels fixed halog entry numbering
  • 221769 12/18/07 06:41 johna left q1 crash
  • 138 12/18/07 06:22 Yoomin Spot again
  • 135 12/18/07 06:14 Kiadtisak right arm coda reboot
  • 115 12/18/07 05:06 johna Major DAQ problems
  • 114 12/18/07 05:04 johna Setting HRS-L to 1.67475 GeV
  • 109 12/18/07 04:38 johna Minor DAQ problem
  • 108 12/18/07 04:26 johna Changing HRS-R to 1.67475
  • 81 12/18/07 02:46 johna HRS-Left to 1.79623 GeV
  • 62 12/18/07 01:37 johna BTA software problems
  • 55 12/18/07 01:10 johna BTA software problem
  • 52 12/18/07 01:01 johna Changing momentum settings
  • 49 12/18/07 00:55 Yoomin spot
  • 42 12/18/07 00:35 Yoomin Target screen
  • 30 12/18/07 00:04 Makis P Shift Summary
  • 28 12/18/07 00:01 minakat target moved to H2 loop3 + target charts
  • 26 12/18/07 00:00 Makis P Current Monotor
  • Monday

  • 25 12/17/07 23:57 Makis P Hall A Tools
  • 20 12/17/07 23:49 minakat target moved to dummy 10cm
  • 15 12/17/07 23:38 minakat target moved to C12 left
  • 12 12/17/07 23:26 Makis P OFFLINE ANALYSIS NOTE
  • 5 12/17/07 23:10 minakat target moved to Fe Left
  • 221765 12/17/07 22:41 minakat target moved to dummy
  • 221760 12/17/07 22:26 minakat target move: H2 loop3
  • 221755 12/17/07 22:14 minakat target move: Pb loop3
  • 221750 12/17/07 22:01 minakat target move from Pb loop2 to He loop2
  • 221741 12/17/07 21:34 minakat target charts at beginning of point #2
  • 221740 12/17/07 21:24 Makis P Set Left HRS at 1.9964 Gev
  • 221735 12/17/07 21:07 minakat target moved to Fe left
  • 221730 12/17/07 20:52 minakat target moved to C12 left
  • 221727 12/17/07 20:45 minakat target moved to Dummy 10cm
  • 221720 12/17/07 20:24 minakat target moved to H2 Loop3
  • 221715 12/17/07 20:12 minakat target moved to Pb loop3
  • 221710 12/17/07 19:58 minakat target moved to He loop2
  • 221705 12/17/07 19:44 minakat target moved to Pb loop2
  • 221702 12/17/07 19:27 Makis P Final raster size
  • 221695 12/17/07 19:18 Makis P Raster size MCC: 3.5x2.8 (2x2mm2 actual)
  • 221682 12/17/07 19:06 Makis P Harp Scan
  • 221681 12/17/07 19:03 Makis P Got beam 5uA, resume data taking
  • 221680 12/17/07 19:03 minakat target moved to He Loop1
  • 221679 12/17/07 17:56 jpchen Loop 2 is operational.
  • 221678 12/17/07 16:42 Makis P RF problem 10 mins estimate
  • 221677 12/17/07 16:29 Makis p MCC wants to do test for uA delivery
  • 221676 12/17/07 16:28 minakat moved target to empty
  • 221666 12/17/07 15:58 Xinhu Yan Day shift Summary
  • 221655 12/17/07 14:54 Ramesh Hall A tools
  • 221652 12/17/07 13:48 Xinhu Yan Harp Scan
  • 221651 12/17/07 13:08 Ramesh Owl shift (Dec 17) summary
  • 221650 12/17/07 13:06 Ramesh Swing shift (Dec 16) summary
  • 221649 12/17/07 13:03 jpchen Target: Cooled down loop 2 last night. Now reduced loop 2 coolant to allwo loop 1 back to normal
  • 221643 12/17/07 11:08 R. Michaels DAQ back up
  • 221641 12/17/07 11:05 atrig Trigger-Download
  • 221638 12/17/07 10:58 Seonho Choi Test of Halog
  • Sunday

  • 221635 12/16/07 21:15 adaq compton stripchart
  • 221608 12/16/07 20:23 minakat Hall A tools
  • 221607 12/16/07 20:19 minakat boiling test c12 reference scan
  • 221602 12/16/07 19:44 Makis P Carbon (left) tgt in
  • 221601 12/16/07 19:41 MinaKat Beam current monitor
  • 221580 12/16/07 17:52 Makis P TGT to Loop 3 H2
  • 221577 12/16/07 17:41 minakat MCC will do spot move
  • 221569 12/16/07 17:11 Makis P TGT Coolant/ESR
  • 221568 12/16/07 17:09 Makis P Loop Temps Charts
  • 221567 12/16/07 17:06 Makis P ATARG Charts
  • 221566 12/16/07 17:06 minakat spot++ 2x2
  • 221549 12/16/07 16:04 Xinhu Yan Day shift Summary
  • 221546 12/16/07 15:53 Ramesh trippy beam again
  • 221497 12/16/07 13:53 xinhu Yan Change Raster size
  • 221478 12/16/07 12:51 Xinhu Yan NMR of right arm is okay now
  • 221477 12/16/07 12:44 folts (techs Right dipole NMR replaced
  • 221476 12/16/07 12:43 Ramesh Controlled access finished
  • 221475 12/16/07 11:26 Ramesh Asked for controlled access
  • 221471 12/16/07 11:15 Y. Zhang target chart
  • 221466 12/16/07 10:59 Ramesh Trippy beam
  • 221460 12/16/07 10:36 Ramesh Error during iocha16 reboot
  • 221441 12/16/07 09:08 Xinhu Yan compton/OTR, new compton run started
  • 221436 12/16/07 08:57 Ramesh Right hrs NMR problem persists
  • 221433 12/16/07 08:44 Xinhu Yan Spot size
  • 221430 12/16/07 08:40 Xinhu Yan Right NMR alarms
  • 221427 12/16/07 08:27 Xinhu Yan Compton and OTR
  • 221426 12/16/07 08:26 Xinhu Yan Ozone chart
  • 221423 12/16/07 08:23 Xinhu Yan Hall A Genral Tools
  • 221419 12/16/07 08:08 Yoomin Owl Shift Summary
  • 221416 12/16/07 08:04 johna Hall A Tools screen for 45 elastic kinematics
  • 221415 12/16/07 08:01 johna shift summary
  • 221412 12/16/07 07:55 johna minor right arm NMR problems
  • 221391 12/16/07 06:09 johna Hall A Tools snapshot for new setting.
  • 221390 12/16/07 06:08 johna Both arms to 35 degrees, 725.96 MeV/c
  • 221383 12/16/07 05:27 Yoomin Target chart
  • 221380 12/16/07 05:14 Yoomin Wrong run number
  • 221377 12/16/07 05:09 Yoomin Correction for automatic entries
  • 221368 12/16/07 04:48 Yoomin Energy measurement start
  • 221362 12/16/07 04:13 Yoomin NaI HV change
  • 221356 12/16/07 04:00 Yoomin NaI HV change
  • 221353 12/16/07 03:54 Yoomin NaI HV change
  • 221347 12/16/07 03:29 Yoomin NaI HV change
  • 221344 12/16/07 03:20 Yoomin NaI HV change
  • 221341 12/16/07 03:12 Yoomin NaI HV change
  • 221338 12/16/07 03:04 Yoomin NaI HV change
  • 221335 12/16/07 02:56 Yoomin NaI HV change
  • 221332 12/16/07 02:47 Yoomin NaI HV change
  • 221329 12/16/07 02:29 Yoomin NaI HV change
  • 221326 12/16/07 02:20 Yoomin NaI HV changed
  • 221323 12/16/07 02:14 Yoomin NaI HV change
  • 221320 12/16/07 02:07 Yoomin NaI HV change
  • 221318 12/16/07 02:06 johna Snapshot: Radiation/Ozone and BPM
  • 221317 12/16/07 02:04 johna Snapshot: Tools and BCM screen
  • 221315 12/16/07 02:00 Yoomin NaI HV change
  • 221312 12/16/07 01:55 johna rotating HRS-R to 15 degrees
  • 221311 12/16/07 01:54 Yoomin NaI HV change
  • 221308 12/16/07 01:45 johna Setting right arm to 0.73647 (cycled Q2/Q3)
  • 221300 12/16/07 01:06 Yoomin NaI HV change
  • 221297 12/16/07 00:58 Yoomin NaI HV change
  • 221291 12/16/07 00:49 Yoomin NaI HV changed
  • 221286 12/16/07 00:39 Yoomin NaI HV change
  • 221283 12/16/07 00:32 Yoomin NaI HV change
  • 221281 12/16/07 00:27 Yoomin NaI HV change
  • 221277 12/16/07 00:03 A. Camsonne NaI pedestal in database updated
  • 221275 12/16/07 00:00 Makis P Current Monitor Chart
  • 221274 12/16/07 00:00 Makis Petrat Shift Summary
  • Saturday

  • 221271 12/15/07 23:46 minakat target screens
  • 221270 12/15/07 23:41 Makis P Hall A Tools
  • 221267 12/15/07 23:29 Nikos sieve slit IN at Left HRS
  • 221266 12/15/07 23:26 lixm NaI HV change
  • 221265 12/15/07 23:25 Nikos setting L/HRS to -2 %
  • 221263 12/15/07 23:22 Nikos R/HRS sieve slit IN
  • 221260 12/15/07 23:15 A. Camsonne Left arm CODA restarted
  • 221259 12/15/07 23:12 Nikos Right HRS at -2% , optics target+open collimator
  • 221256 12/15/07 23:09 lixm NaI HV change
  • 221253 12/15/07 23:00 lixm NaI HV change
  • 221250 12/15/07 22:52 A. Camsonne reprocess run 3362 to 3763
  • 221248 12/15/07 22:52 lixm NaI HV change
  • 221246 12/15/07 22:41 lixm NaI HV change
  • 221244 12/15/07 22:28 Nikos HV scans for NaI
  • 221242 12/15/07 22:21 Nikos R/HRS to -2 %
  • 221241 12/15/07 22:19 Adaq NaI HV change
  • 221240 12/15/07 22:17 Nikos moving to optics target
  • 221235 12/15/07 22:03 Adaq Ozone/Rad Mon
  • 221224 12/15/07 21:11 Makis P Offline Analysis Note for Elastic Runs
  • 221205 12/15/07 20:05 minakat Target moved to dummy
  • 221198 12/15/07 19:08 A. Camsonne Left Arm set to 735.14 MeV/c Sieve slit in
  • 221197 12/15/07 19:04 A. Camsonne NaI HV restored to 712 MeV/c HV
  • 221192 12/15/07 18:56 A. Camsonne NaI HV +100v for run 4056
  • 221189 12/15/07 18:49 Makis P Beam is back 16:45
  • 221187 12/15/07 18:38 A. Camsonne NaI HV change +50 V
  • 221186 12/15/07 18:35 Makis P RF Problem Down for 15 mins
  • 221183 12/15/07 18:26 A. Camsonne NaI HV change -50 V run 4054
  • 221181 12/15/07 18:22 A. Camsonne NaI HV -100V for run 4053
  • 221177 12/15/07 18:11 A. Camsonne NaI HV Change for run 4052
  • 221176 12/15/07 18:09 A. Camsonne HV scan at 4%
  • 221169 12/15/07 17:19 A. Camsonne NaI HV restored to 712 MeV setting for +2%
  • 221164 12/15/07 17:10 A. Camsonne NaI HV set to 2000V
  • 221160 12/15/07 17:00 A. Camsonne NaI HV change
  • 221155 12/15/07 16:51 A. Camsonne NaI HV Change
  • 221151 12/15/07 16:43 A. Camsonne NaI HV change
  • 221149 12/15/07 16:37 A. Camsonne NaI HV change for run 4045
  • 221147 12/15/07 16:35 A. Camsonne NaI HV change
  • 221141 12/15/07 16:14 minakat target charts
  • 221140 12/15/07 16:08 Lindgren Shift Summary
  • 221138 12/15/07 16:00 lindgren rad mon,bcm,bpm
  • 221137 12/15/07 15:58 lindgren hall a tools
  • 221135 12/15/07 15:49 Aniol spot++
  • 221133 12/15/07 15:22 Ramesh Both hrs's at 23 deg, sieve slits of both on
  • 221125 12/15/07 13:59 A. Camsonne background
  • 221124 12/15/07 13:58 A. Camsonne background
  • 221115 12/15/07 12:38 R.A. Lindgre Mid shift summary
  • 221113 12/15/07 12:25 R.A. Lindgre Production running
  • 221110 12/15/07 12:11 Aniol S2 tdc double peaks in 4034
  • 221104 12/15/07 11:22 R.A.Lindgren Right HRS in Motion
  • 221097 12/15/07 10:12 R.A. Lindgre Momentum change
  • 221085 12/15/07 09:12 R.A.Lindgren Day Shift Starting
  • 221071 12/15/07 08:04 x. jiang owl shift summary
  • 221043 12/15/07 06:31 x. jiang magnet stripcharts
  • 221042 12/15/07 06:30 X. Jiang stripcharts
  • 221041 12/15/07 06:29 X. Jiang strip charts
  • 220973 12/15/07 02:01 X. Jiang tools screen
  • 220945 12/15/07 00:00 Seonho Choi Shift summary (Owl 12/14/2007)
  • Friday

  • 220902 12/14/07 21:32 Seonho Choi BPM/BCM
  • 220901 12/14/07 21:32 Seonho Choi Compton
  • 220900 12/14/07 21:31 Seonho Choi Ozone/Radiation
  • 220899 12/14/07 21:31 Seonho Choi Hall-A Tools
  • 220844 12/14/07 18:20 lixm target screen
  • 220837 12/14/07 17:58 lixm target screen
  • 220836 12/14/07 17:57 Seonho Choi raster after the beam back
  • 220829 12/14/07 17:44 A. Camsonne Compton issue
  • 220828 12/14/07 17:44 Seonho Choi Compton Back
  • 220825 12/14/07 17:20 Seonho Choi Compton DAQ/computer failure
  • 220797 12/14/07 13:42 averett radiation strip chart
  • 220796 12/14/07 13:41 averett compton, OTR stripcharts
  • 220795 12/14/07 13:40 averett BPM/BCM strip charts
  • 220794 12/14/07 13:38 averett HA tools
  • 220752 12/14/07 10:34 Lindgren Target mislabeled in halog
  • 220729 12/14/07 08:47 averett HV alarm
  • 220725 12/14/07 08:35 Aniol r3928 missing adcs
  • 220718 12/14/07 08:02 X. Jiang Owl shift summary
  • 220717 12/14/07 08:00 Liyang Jiang Target charts
  • 220700 12/14/07 07:02 X. Jiang Tools screen
  • 220675 12/14/07 05:25 Liyang Jiang Target theta between arms
  • 220670 12/14/07 05:15 Liyang Jiang Target charts
  • 220657 12/14/07 04:43 X. jiang Tools screen
  • 220652 12/14/07 04:24 Yoomin Compton reboot.
  • 220650 12/14/07 04:13 Liyang Jiang Target charts
  • 220630 12/14/07 03:08 Liyang Jiang Target charts
  • 220625 12/14/07 02:49 X. Jiang Run 3902, He target, PL=0.64259
  • 220623 12/14/07 02:45 Yoomin NaI HV change
  • 220606 12/14/07 01:53 Liyang Jiang target
  • 220600 12/14/07 01:40 X. Jiang Tools screen
  • 220597 12/14/07 01:30 X. Jiang Tools screen
  • 220593 12/14/07 01:09 X. Jiang ight arm momentum takes a long time to settle.
  • 220591 12/14/07 01:06 Liyang Jiang He target left arm Momemtum 665.085
  • 220587 12/14/07 00:47 Liyang Jiang Target Y of He target
  • 220579 12/14/07 00:27 Liyang Jiang Empty histogram of poin rejector1 adc 2
  • 220574 12/14/07 00:15 X. Jiang Tools screen and beam current
  • 220573 12/14/07 00:14 x. jiang raster from spot_l
  • 220570 12/14/07 00:02 Seonho Choi Shift summary (2/2) for 60 degrees
  • Thursday

  • 220564 12/13/07 23:54 Chiranjib TARGET----->>> C
  • 220558 12/13/07 23:42 Seonho Choi Some strange thing as regard to HV
  • 220556 12/13/07 23:36 Chiranjib TARGET--->>> Fe
  • 220553 12/13/07 23:27 Chiranjib TARGET--->> C
  • 220550 12/13/07 23:17 Chiranjib TARGET----->>> DUMMY
  • 220547 12/13/07 23:05 Chiranjib TARGET------>>> Pb
  • 220544 12/13/07 22:53 Seonho Choi Configuration Change
  • 220543 12/13/07 22:29 Chiranjib TARGET--->>> He4
  • 220542 12/13/07 22:16 Seonho Choi 739MeV/90 deg done. Moving to 60 degrees.
  • 220537 12/13/07 22:00 Seonho Choi Shift summary for 739MeV, 90 degrees
  • 220536 12/13/07 21:59 Chiranjib - TARGET--->>> Pb
  • 220527 12/13/07 21:39 Chiranjib TARGET------>>>H2
  • 220524 12/13/07 21:25 Seonho Choi Hall-A Tools
  • 220519 12/13/07 20:59 Chiranjib TARGET ---->>Pb
  • 220516 12/13/07 20:50 chiranjib TARGET---->> H2
  • 220513 12/13/07 20:30 Chiranjib TARTET ----->>> C
  • 220511 12/13/07 20:25 Seonho Choi BPM/BCM
  • 220510 12/13/07 20:25 Seonho Choi Ozone/Radiation
  • 220509 12/13/07 20:24 Seonho Choi Compton rates
  • 220505 12/13/07 19:52 Chiranjib TARGET---->> Fe
  • 220502 12/13/07 19:43 Chiranjib TARGET SCREENS II
  • 220501 12/13/07 19:43 Chiranjib TARET SCREENS I
  • 220498 12/13/07 19:35 Chiranjib TARGET ----->>> C
  • 220493 12/13/07 19:23 Chiranjib TARGET ----->> H2
  • 220488 12/13/07 19:05 chiranjib TARGET---->> Pb
  • 220483 12/13/07 18:42 Chiranjib TARGET----->>> Fe
  • 220480 12/13/07 18:25 Chiranjib TARGET ---->> C
  • 220477 12/13/07 18:13 Chiranjib TARGET----->>> H2
  • 220474 12/13/07 17:54 Chiranjib TARGET----->> Pb
  • 220469 12/13/07 17:38 Chiranjib TARGET --->> He 4
  • 220466 12/13/07 17:28 Chiranjib TARGET ----->>>Pb
  • 220459 12/13/07 17:02 Chiranjib TARGET--->> H2
  • 220450 12/13/07 16:50 Chiranjib TARGET ---->>> DUMMY
  • 220449 12/13/07 16:49 Seonho Choi Raster Check
  • 220444 12/13/07 16:22 Seonho Choi Beam back
  • 220443 12/13/07 16:19 Chiranjib TAREGT --->> Fe
  • 220440 12/13/07 16:11 folts Spectrometer move and polatity change
  • 220438 12/13/07 16:05 Lindgren Day Shift Summary
  • 220435 12/13/07 15:39 Aniol target pages
  • 220434 12/13/07 15:34 Lindgren BCM/BPM
  • 220433 12/13/07 15:33 Lindgren Ozone strip chart
  • 220432 12/13/07 15:32 Lindgren Hall A Tools
  • 220431 12/13/07 12:40 Lindgren Mid-shift Summary
  • 220417 12/13/07 11:41 Lindgren Production running
  • 220413 12/13/07 11:02 Lindgren Recycle Q2/Q3
  • 220412 12/13/07 11:00 Xinhu Yan Target Status
  • 220393 12/13/07 09:11 Lindgren 8:00 - 9:00 Production Running
  • 220383 12/13/07 08:06 jpchen owl shift summary
  • 220378 12/13/07 07:31 jpchen left to 175.47 MeV
  • 220375 12/13/07 07:07 jpchen change right p to 121.169 MeV
  • 220359 12/13/07 06:06 jpchen left to 188.198 MeV
  • 220340 12/13/07 05:00 jpchen left p to 201.85 MeV
  • 220326 12/13/07 03:47 jpchen left to 216.492 MeV
  • 220321 12/13/07 03:23 jpchen target screens
  • 220320 12/13/07 03:22 jpchen HVs: Left and Beamline
  • 220319 12/13/07 03:20 jpchen HVs: right top/bottom
  • 220318 12/13/07 03:15 jpchen Radiation/Ozone
  • 220317 12/13/07 03:15 jpchen BPM/BCM
  • 220316 12/13/07 03:14 jpchen Compton/OTR
  • 220315 12/13/07 03:13 jpchen Hall A general tools screen
  • 220310 12/13/07 03:01 jpchen Raster size changed (by itself!) during 12/11 swing
  • 220297 12/13/07 01:49 jpchen left to 232.196 MeV
  • 220292 12/13/07 01:25 jpchen change right to 148.021 MeV
  • 220276 12/13/07 00:02 Xiaodong Jia swing shift summary, good event #
  • Wednesday

  • 220275 12/12/07 23:48 X. Jiang Run 3817, Fe target, Y_tar is OK
  • 220272 12/12/07 23:40 X. Jiang run 3812, Pb target, Ytar
  • 220271 12/12/07 23:27 X. Jiang Run 3811, H2 target, Yar.
  • 220270 12/12/07 23:25 O. Glamazdin Shif Summary for 12/11/07 Swing Shift
  • 220269 12/12/07 22:58 X. Jiang run 21818, Pb target. Ytar
  • 220268 12/12/07 22:45 X. Jiang Pb target, run 3805. Ytar
  • 220257 12/12/07 21:42 X. Jiang Tools screen
  • 220242 12/12/07 20:34 X. Jiang Tools screen
  • 220239 12/12/07 20:06 mcnulty Target Screens
  • 220238 12/12/07 20:04 X. Jiang Compton chart and OTR
  • 220223 12/12/07 18:39 X. Jiang Tools screen
  • 220220 12/12/07 18:23 X. jiang raster
  • 220207 12/12/07 17:13 X. Jiang Hall A tool screens shot
  • 220197 12/12/07 15:57 Xinhu Yan Day shift summary
  • 220192 12/12/07 15:47 jpchen Raster size now 2x2
  • 220187 12/12/07 15:42 jpchen raster size too large in x
  • 220184 12/12/07 15:34 Adaq run 3793 p should 0.17547 MeV
  • 220162 12/12/07 12:37 Xinhu Yan Pion Rejector Pedstal changed
  • 220158 12/12/07 11:48 Xinhu Yan Momentum changed
  • 220151 12/12/07 11:29 Xinhu Yan Pion Rejector Pedstal changed
  • 220142 12/12/07 10:27 R. Michaels scaler history
  • 220139 12/12/07 09:59 R. Michaels re: u10, d10
  • 220138 12/12/07 09:42 Xinhu Yan Beam Off
  • 220137 12/12/07 09:32 Seonho Choi BCM u10/d10 flaky
  • 220134 12/12/07 09:07 Xinhu Yan Spot++
  • 220133 12/12/07 08:59 Xinhu Yan Target Charts
  • 220132 12/12/07 08:45 Xinhu Yan Radiation Ozone status
  • 220131 12/12/07 08:43 Xinhu Yan Compton and OTR charts
  • 220130 12/12/07 08:42 Xinhu Yan BPM and BCM Charts
  • 220129 12/12/07 08:41 Xinhu Yan Hall A General Tool
  • 220123 12/12/07 08:10 Y. Zhang Owl Shift Summary
  • 220108 12/12/07 05:10 lixm Target screen
  • 220107 12/12/07 05:07 Y. Zhang Compton & OTR
  • 220106 12/12/07 05:06 Y. Zhang Radiation_ozone
  • 220103 12/12/07 05:05 Y. Zhang BCM&BPM
  • 220102 12/12/07 05:03 Y. Zhang tools page
  • Tuesday

  • 220052 12/11/07 23:15 O. Glamazdin Right arm dead-time
  • 220048 12/11/07 23:04 A. Camsonne Right arm dead-time
  • 220042 12/11/07 22:40 jpchen Deadtime in end run for right DAQ
  • 220013 12/11/07 19:52 mcnulty Target Screens
  • 220012 12/11/07 19:46 O. Glamazdin BCM and BPM stripcharts
  • 220011 12/11/07 19:44 O. Glamazdin Compton and OTR stripcharts
  • 220010 12/11/07 19:42 O. Glamazdin Ozon Stripcharts
  • 220009 12/11/07 19:39 O. Glamazdin Hall A Tools Page
  • 219996 12/11/07 18:10 O. Glamazdin Compton DAQ reboot
  • 219987 12/11/07 17:10 A. Camsonne NaI HV change
  • 219975 12/11/07 15:58 Aniol shift summary
  • 219956 12/11/07 14:25 Aniol set P0L to 0.307261
  • 219953 12/11/07 13:46 Aniol P0R set to 0.353454
  • 219950 12/11/07 13:30 A. Camsonne NaI status
  • 219949 12/11/07 13:29 A. Camsonne NaI status
  • 219948 12/11/07 13:25 Aniol alarm handler on HV not working?
  • 219943 12/11/07 12:55 Aniol tgt shots
  • 219942 12/11/07 12:52 Aniol ozone, rad monitors
  • 219941 12/11/07 12:50 Aniol bcm,bpm
  • 219940 12/11/07 12:47 Aniol tools, bcm
  • 219927 12/11/07 12:00 Aniol raster size
  • 219922 12/11/07 11:25 Aniol run 3735,P0L=0.329549
  • 219919 12/11/07 11:12 Aniol P0L set to 0.329549
  • 219908 12/11/07 09:49 Aniol change P0R to 0.379093
  • 219905 12/11/07 09:20 Aniol Pb tgt in at 40 uA
  • 219898 12/11/07 08:48 Aniol change P0L to 0.353454
  • 219890 12/11/07 08:07 Jose Udias Owl shift summary
  • 219873 12/11/07 06:28 R. Michaels R-HRS scaler checkout
  • 219870 12/11/07 06:21 R. Michaels L-HRS scaler checkout
  • 219861 12/11/07 05:43 R. Michaels target charts
  • 219860 12/11/07 05:42 R. Michaels target screens
  • 219858 12/11/07 05:36 jose udias BPM/BCM strip charts
  • 219854 12/11/07 05:17 Jose Udias spot image
  • 219849 12/11/07 04:50 Jose Udias Compton / OTR strips
  • 219848 12/11/07 04:48 jose udias Radiation / Ozone strips
  • 219847 12/11/07 04:46 Jose Udias Hall A tools
  • 219808 12/11/07 01:32 R. Michaels reboot iocha16
  • 219784 12/11/07 00:02 Ole Hansen Swing shift summary
  • Monday

  • 219777 12/10/07 23:27 Ole Hansen HRS-R to 538 MeV
  • 219776 12/10/07 23:27 Ole Hansen HRS-L to 577 MeV
  • 219761 12/10/07 22:54 Ole Hansen Target screenshot #2
  • 219760 12/10/07 22:52 Ole Hansen Target screenshots
  • 219757 12/10/07 22:49 Ole Hansen BCM/BPM strip charts
  • 219756 12/10/07 22:48 Ole Hansen O3 Rad stripcharts
  • 219751 12/10/07 22:38 Ole Hansen Use RIGHT foil targets
  • 219742 12/10/07 22:19 Ole Hansen HRS-L to 618.9 MeV
  • 219737 12/10/07 21:58 Ole Hansen Target vaccum
  • 219736 12/10/07 21:55 Ole Hansen Loop3 excursion
  • 219733 12/10/07 21:42 Ole Hansen HRS-R to 618.9 MeV
  • 219732 12/10/07 21:42 Ole Hansen HRS-L to 663.8 MeV
  • 219725 12/10/07 21:24 Ole Hansen Hall A Tools
  • 219722 12/10/07 21:21 Ole Hansen spot results
  • 219712 12/10/07 20:56 Ole Hansen Target to Fe
  • 219711 12/10/07 20:56 Ole Hansen Harp results
  • 219710 12/10/07 20:55 Ole Hansen Compton rate
  • 219709 12/10/07 20:53 Ole Hansen OTR good
  • 219708 12/10/07 20:31 Ole Hansen Beam returns
  • 219706 12/10/07 18:47 A. Camsonne NaI HV change
  • 219705 12/10/07 18:08 Yoomin BCM d3, u10 for right arm
  • 219704 12/10/07 16:28 Techs Spectometer angle / Polarity
  • 219702 12/10/07 16:02 LeRose Shift Summary
  • 219701 12/10/07 14:58 Aniol tgt controls
  • 219700 12/10/07 14:57 Aniol tgt charts
  • 219699 12/10/07 14:52 LeRose BCM/BPM
  • 219698 12/10/07 14:37 LeRose Ozone/Rad Mon
  • 219697 12/10/07 14:17 LeRose Controlled Access (left arm polarity change) & move to 120°
  • 219674 12/10/07 13:06 LeRose New Left P0 (0.11083)
  • 219669 12/10/07 12:46 LeRose Left P0=0.12249 GeV/c
  • 219639 12/10/07 10:51 LeRose New Left Arm P0
  • 219637 12/10/07 10:38 aniol spot++
  • 219622 12/10/07 09:40 LeRose Beam Back
  • 219619 12/10/07 09:36 LeRose Hall A Tools
  • 219618 12/10/07 08:49 LeRose Controlled Access (left arm polarity change)
  • 219609 12/10/07 08:00 Holmstrom Shift Summary- Owl
  • 219606 12/10/07 07:45 Holmstrom Change Momentums
  • 219594 12/10/07 06:58 Holmstrom BCM/BPM
  • 219593 12/10/07 06:58 Holmstrom Ozone/Rad Mon.
  • 219592 12/10/07 06:57 Holmstrom hall A Tools
  • 219591 12/10/07 06:54 Holmstrom We briefly lost T1+T2
  • 219584 12/10/07 06:27 Holmstrom Momentum Change
  • 219569 12/10/07 05:04 Holmstrom Momentum changes
  • 219544 12/10/07 03:36 Holmstrom Left Arm Momentum Change
  • 219529 12/10/07 02:20 Holmstrom Change Momentums
  • 219504 12/10/07 00:42 Holmstrom Momentum Change
  • 219503 12/10/07 00:41 A. Camsonne NaI HV increased
  • 219495 12/10/07 00:02 C. Fernandez Swing shift summary
  • Sunday

  • 219438 12/09/07 20:21 Vignote Target snapshot
  • 219435 12/09/07 20:16 C. Fernandez Compton snapshot
  • 219434 12/09/07 20:15 C. Fernandez BPM/BCN snapshot
  • 219433 12/09/07 20:15 C. Fernandez Current snapshot
  • 219432 12/09/07 20:14 C. Fernandez Hall A tools snapshot
  • 219418 12/09/07 19:34 Spiegel/Tech Magnet liquid levels
  • 219412 12/09/07 19:10 C. Fernandez Return to beam permit
  • 219411 12/09/07 19:04 Spiegel/Tech LEFT Dipole
  • 219410 12/09/07 17:32 C. Fernandez Controlled access
  • 219409 12/09/07 16:43 C. Fernandez Helium flow status
  • 219406 12/09/07 16:00 Seonho Choi Update on He level crisis in L-dipole
  • 219404 12/09/07 15:55 Seonho Choi Shift summary (runs taken)
  • 219402 12/09/07 15:44 Seonho Choi Hall-A Tools
  • 219399 12/09/07 15:30 Seonho Choi Wait and see that He level for L dipole come back to normal
  • 219395 12/09/07 15:13 Seonho Choi He level low for L-HRS dipole
  • 219394 12/09/07 15:12 Seonho Choi L-HRS dipole trip
  • 219375 12/09/07 14:05 yez Target Screen Grab,Dec.9th
  • 219374 12/09/07 14:03 yez Target Screen Grab,Dec.9th
  • 219367 12/09/07 13:42 Seonho Choi Compton
  • 219366 12/09/07 13:42 Seonho Choi Compton
  • 219365 12/09/07 13:41 Seonho Choi BPM/BCM
  • 219364 12/09/07 13:41 Seonho Choi Ozone/Radiation
  • 219363 12/09/07 13:40 Seonho Choi Hall-A Tools
  • 219358 12/09/07 13:17 Seonho Choi Raster
  • 219350 12/09/07 12:58 Seonho Choi Beam back finally
  • 219348 12/09/07 12:56 Seonho Choi Post-Halog Entry: Helium Flow shortage/recovery for magnets
  • 219347 12/09/07 12:27 Seonho Choi Beam off again
  • 219342 12/09/07 12:14 Seonho Choi Beam back
  • 219339 12/09/07 10:45 Seonho Choi MCC took away the beam
  • 219330 12/09/07 10:07 Seonho Choi MCC to work on Hall C beam problem
  • 219311 12/09/07 09:05 Seonho Choi DAQ time cut down
  • 219300 12/09/07 08:07 jose owl shift summary
  • 219281 12/09/07 06:56 jose udias Restarted CODA in LEFT ARM
  • 219255 12/09/07 05:18 jose udias hall A tools
  • 219254 12/09/07 05:17 jose udias BPM strip chart
  • 219252 12/09/07 05:17 jose udias ozone/rad./BCM stripcharts
  • 219197 12/09/07 02:27 jose Hall A tools snapshot
  • 219191 12/09/07 02:08 Yoomin target chart
  • 219183 12/09/07 01:30 jose raster size
  • 219182 12/09/07 01:29 Yoomin Target Empty>>Fe right
  • 219175 12/09/07 01:07 Changing in CHL flow
  • 219172 12/09/07 00:00 C. Fernandez Swing shift summary
  • Saturday

  • 219170 12/08/07 23:17 C. Fernandez Tools snapshot
  • 219169 12/08/07 22:23 C. Fernandez Tools snapshot
  • 219168 12/08/07 21:45 C. Fernandez Tools snapshot
  • 219161 12/08/07 21:13 C. Fernandez Tool snapshot
  • 219148 12/08/07 19:02 C. Fernandez Q3 failure in L and R magnets
  • 219147 12/08/07 18:44 C. Fernandez Left arm Q3 problem
  • 219146 12/08/07 18:05 A. Camsonne Helium level low on magnets
  • 219145 12/08/07 18:04 C. Fernandez Right arm Q3
  • 219144 12/08/07 18:04 C. Fernandez Magnet Cryogen Levels
  • 219143 12/08/07 16:51 A. Camsonne NaI HV increased by 80 V
  • 219139 12/08/07 16:01 J.L.Herraiz Day Shift Summary
  • 219134 12/08/07 09:46 J.L.Herraiz Going to Controlled Access
  • 219133 12/08/07 08:10 jpchen lowered CHL flow
  • 219132 12/08/07 08:03 jpchen owl shift summary
  • 219128 12/08/07 06:43 jpchen both to 90 degrees
  • 219127 12/08/07 06:05 jpchen change both to 712.65 MeV
  • 219126 12/08/07 05:52 jpchen Moved target to EMPTY
  • 219125 12/08/07 05:31 jpchen BCM chart for Pb+LH2 boiling test
  • 219124 12/08/07 05:30 jpchen completed data taking at 1.256 GeV, energy change started
  • 219121 12/08/07 05:28 jpchen Pb(+LH2) vs current test
  • 219104 12/08/07 04:52 jpchen deadtime high for 3527/3528
  • 219103 12/08/07 04:50 jpchen TS0 frozen, reboot/restarted coda
  • 219092 12/08/07 04:20 jpchen runs 3516-3524 are at 20 degrees
  • 219091 12/08/07 04:15 jpchen Left to 33 degree, 1034.956 MeV
  • 219089 12/08/07 04:09 jpchen He boiling test completed
  • 219062 12/08/07 03:00 lixm target screens
  • 219055 12/08/07 02:47 jpchen BPM/BCM
  • 219054 12/08/07 02:45 jpchen Radiation/Ozone
  • 219053 12/08/07 02:44 jpchen Compton/OTR
  • 219052 12/08/07 02:43 jpchen Left to 1233.873 MeV, Tools screen
  • 219049 12/08/07 02:28 jpchen left to 20 degree
  • 219041 12/08/07 02:08 jpchen moved left sieve out
  • 219035 12/08/07 01:46 jpchen moved left sieve in and to optics target
  • 219034 12/08/07 01:46 jpchen boiling test runs with 12C
  • 219003 12/08/07 00:01 C. Fernandez Swing shift summary
  • Friday

  • 218957 12/07/07 20:32 C. Fernandez Tools snapshot
  • 218950 12/07/07 20:16 C. Fernandez Current snapshot
  • 218946 12/07/07 20:09 C. Fernandez BPM/BCN snapshot
  • 218945 12/07/07 20:09 C. Fernandez Compton snapshot
  • 218924 12/07/07 19:18 Vignote Target snapshot
  • 218901 12/07/07 18:12 A. Camsonne NaI HV
  • 218863 12/07/07 16:05 Xinhu Yan Day shift Summary
  • 218852 12/07/07 15:23 Xinhu Yan Recycle left arm Q2/Q3
  • 218843 12/07/07 14:54 A. Camsonne NaI HV change
  • 218834 12/07/07 14:27 A. Camsonne NaI HV change
  • 218833 12/07/07 14:26 J.L. Herraiz Target charts
  • 218824 12/07/07 14:02 Xinhu Yan Beam back
  • 218823 12/07/07 14:01 A. Camsonne Lowering left arm momentum for NaI check
  • 218821 12/07/07 13:48 Spiegel/Tech Diffuser Cooler FSD fault
  • 218819 12/07/07 12:51 A. Camsonne NaI HV change
  • 218818 12/07/07 12:48 Xinhu Yan beam away
  • 218809 12/07/07 11:48 A. Camsonne NaI HV reduced bottom part
  • 218797 12/07/07 11:07 Xinhu Yan OTR and Compton Charts
  • 218795 12/07/07 11:06 Xinhu Yan Bcm and BpM charts
  • 218788 12/07/07 10:37 Xinhu Yan Diffuser Water flow fixed
  • 218787 12/07/07 09:09 Xinhu Yan Hall A Control Tool
  • 218784 12/07/07 08:49 Xinhu Yan Water flow diffuser Problem
  • 218783 12/07/07 08:05 Y. Zhang Owl Shift Summary
  • 218781 12/07/07 07:25 Yoomin NaI HV Changed
  • 218763 12/07/07 04:57 Y. Zhang target charts
  • 218761 12/07/07 04:54 Y. Zhang Ozone & Radiation Monitors
  • 218760 12/07/07 04:52 Y. Zhang BPM chart
  • 218759 12/07/07 04:51 Y. Zhang BCM & OTR chart
  • 218757 12/07/07 04:48 Y. Zhang Hall A general tools chart
  • 218739 12/07/07 02:31 Y. Zhang coda right arm
  • 218738 12/07/07 02:28 A. Camsonne Rebooted ROC1
  • 218731 12/07/07 02:15 Y. Zhang left arm momentum change
  • 218729 12/07/07 01:57 Y. Zhang coda reboot
  • 218718 12/07/07 01:02 Y. Zhang Momentum Changing
  • 218710 12/07/07 00:02 C. Fernandez Swing shift summary
  • Thursday

  • 218689 12/06/07 22:41 A. Camsonne NaI HV changed
  • 218681 12/06/07 22:13 C. Fernandez BCM/BPM Snapshot
  • 218680 12/06/07 22:12 Vignote target snapshot
  • 218679 12/06/07 22:12 C. Fernandez Compton snapshot
  • 218678 12/06/07 22:11 C. Fernandez Ozone snapshot
  • 218677 12/06/07 22:10 C. Fernandez Current snapshot
  • 218676 12/06/07 22:08 C. Fernandez tools snapshot
  • 218659 12/06/07 20:59 Jack VDC gas mixing system purged
  • 218646 12/06/07 20:02 Vignote Target snapshot
  • 218627 12/06/07 18:46 Adaq NaI HV changed
  • 218618 12/06/07 18:06 A. Camsonne NaI HV changed
  • 218594 12/06/07 16:03 R. Michaels day shift summary
  • 218591 12/06/07 16:00 J.Lopez Pb target
  • 218589 12/06/07 15:52 R. Michaels spot_L and spot_R
  • 218586 12/06/07 15:42 R. Michaels BPM/BCM screens
  • 218585 12/06/07 15:41 R. Michaels radiation-ozone strip chart
  • 218584 12/06/07 15:40 R. Michaels tools screen
  • 218579 12/06/07 15:30 J.Lopez Target charts
  • 218574 12/06/07 14:39 R. Michaels P0 changes
  • 218567 12/06/07 14:16 J.Lopez Running with 30uA with Pb
  • 218550 12/06/07 13:13 R. Michaels P0L changed
  • 218543 12/06/07 12:21 R. Michaels P0R change
  • 218534 12/06/07 11:47 R. Michaels P0L change
  • 218527 12/06/07 11:25 J.L. Herraiz Spot++
  • 218526 12/06/07 11:22 R. Michaels beam returns.
  • 218523 12/06/07 10:40 A. Camsonne NaI HV increased
  • 218520 12/06/07 10:11 R. Michaels beam returns
  • 218517 12/06/07 10:08 Xinhu Yan Pion rejector 2 calibration
  • 218512 12/06/07 09:47 Xinhu Yan Pion Rejector Calibration
  • 218511 12/06/07 08:49 Xinhu Yan Pion Rejector 1 calibration
  • 218510 12/06/07 08:47 Xinhu Yan Pion Rejector 1 calibration
  • 218505 12/06/07 08:09 R. Michaels P0 changes
  • 218481 12/06/07 07:02 jpchen left to 1053.222 MeV
  • 218473 12/06/07 06:24 jpchen Left to 1129.621 MeV, right to 1053.222 MeV
  • 218466 12/06/07 06:08 jpchen Compton/OTR
  • 218465 12/06/07 06:07 jpchen BPM/BCM
  • 218464 12/06/07 06:06 jpchen radiation/ozone
  • 218463 12/06/07 06:05 jpchen Hall A tools screen
  • 218461 12/06/07 06:04 jpchen raster
  • 218455 12/06/07 05:53 jpchen harp scan 1h03A
  • 218454 12/06/07 05:49 jpchen raster check
  • 218449 12/06/07 05:37 jpchen Compton ~7KHz @ 50 uA
  • 218448 12/06/07 04:48 jpchen target screens
  • 218446 12/06/07 01:17 jpchen tuning beam
  • 218445 12/06/07 00:26 jpchen no beam
  • 218443 12/06/07 00:00 C. Fernandez Swing shift summary
  • Wednesday

  • 218437 12/05/07 21:33 A. Camsonne NaI HV change
  • 218430 12/05/07 18:08 Xinhu Yan calibration for Pion Rejector 1
  • 218429 12/05/07 17:01 C. Fernandez Left Spectrometer
  • 218428 12/05/07 16:35 Jack Left Q2 power supply status
  • 218427 12/05/07 15:58 L. Zhu Shift Summary
  • 218424 12/05/07 13:07 L. Zhu controlled access
  • 218422 12/05/07 08:21 A. Camsonne Lead target plots
  • 218421 12/05/07 08:09 L. Zhu beam pass change
  • 218417 12/05/07 08:01 Z. Meziani Shift Summary
  • 218410 12/05/07 07:23 Meziani Beam position scan on Pb target
  • 218395 12/05/07 06:55 Meziani Summary of Optics runs
  • 218378 12/05/07 05:51 Meziani MOving spectrometers to 40 deg
  • 218377 12/05/07 05:48 A. Camsonne Run 3353
  • 218372 12/05/07 04:59 Meziani Beam SCAN study
  • 218355 12/05/07 02:43 A. Camsonne NaI HV
  • 218352 12/05/07 02:16 A. Camsonne NaI HV change
  • 218339 12/05/07 01:23 A. Camsonne NaI HV change
  • 218332 12/05/07 00:49 A. Camsonne NaI HV change
  • 218331 12/05/07 00:42 jpchen Hydrogen tank pressure transducer problem.
  • 218326 12/05/07 00:27 A. Camsonne reduced HV NaI
  • 218323 12/05/07 00:01 C. Fernandez Swing shift summary
  • Tuesday

  • 218309 12/04/07 22:53 A. Camsonne NaI HV restored to production HV values and pedestal subtraction back on
  • 218302 12/04/07 22:24 C. Fernandez Ozone snapshot
  • 218301 12/04/07 22:24 C. Fernandez Compton
  • 218300 12/04/07 22:23 C. Fernandez Current
  • 218299 12/04/07 22:23 C. Fernandez Tools
  • 218298 12/04/07 22:21 jpchen raster check
  • 218294 12/04/07 21:07 Vignote heat load incremented
  • 218293 12/04/07 21:01 V. Sulkosky LHRS and RHRS Momentum Set
  • 218292 12/04/07 20:15 Jack (Techs) Left Q2 power supply
  • 218291 12/04/07 20:11 Jack Gas system shutdown earlier today
  • 218289 12/04/07 19:11 A. Camsonne Run 3333 cosmics pedestal subtraction off
  • 218286 12/04/07 19:01 A. Camsonne NaI HV changed
  • 218283 12/04/07 17:16 ewing checklist
  • 218282 12/04/07 17:16 A. Camsonne NaI HV lowered for cosmics
  • 218279 12/04/07 16:03 E.Chudakov Day shift summary
  • 218276 12/04/07 13:22 J. Lopez Gas Shed - Negative preassure
  • 218275 12/04/07 10:49 J.Lopez Heat load reduced to 60W
  • 218274 12/04/07 09:06 J. Lopez Reducing the Heat Load in the target
  • 218271 12/04/07 08:00 Meziani Shift Summary
  • 218253 12/04/07 04:20 Meziani HVs left arm
  • 218251 12/04/07 04:18 Meziani HVs Right Arm
  • 218250 12/04/07 04:17 Y. Zhang target chart
  • 218249 12/04/07 04:16 Meziani Compton rates/OTR charts
  • 218248 12/04/07 04:15 Meziani Ozone/Radiation Monitors Charts
  • 218247 12/04/07 04:14 Meziani Hall A tools
  • 218235 12/04/07 01:04 Meziani Changing momentum of right arm
  • 218226 12/04/07 00:03 C. Fernandez Swing shift summary
  • Monday

  • 218223 12/03/07 23:46 Vignote Target snapshot
  • 218221 12/03/07 23:40 C. Fernandez BPM
  • 218220 12/03/07 23:40 C. Fernandez BCM
  • 218219 12/03/07 23:39 C. Fernandez BCN/BPM
  • 218218 12/03/07 23:38 C. Fernandez Ozone snapshot
  • 218217 12/03/07 23:37 C. Fernandez Tools snapshot
  • 218213 12/03/07 21:14 A. Camsonne NaI HV change for cosmics
  • 218212 12/03/07 20:53 Vignote Target heat load incremented
  • 218208 12/03/07 17:15 Spiegel / Te Pre-beam checklist
  • 218207 12/03/07 16:47 jpchen reduced Helium pressure to 170 psi
  • 218206 12/03/07 16:19 A. Camsonne NaI Status
  • 218205 12/03/07 16:04 E.Chudakov Day shift summary
  • 218203 12/03/07 14:29 J.Lopez Target heat load reduced
  • 218202 12/03/07 14:29 Yoomin Compton rebooted.
  • 218193 12/03/07 11:09 E.Chudakov Restricted access
  • 218192 12/03/07 10:45 Shabestari Shift summary, Tuesday 11/27/2007 (Problem with HALOG)
  • 218190 12/03/07 08:21 E.Chudakov Control access request
  • 218186 12/03/07 08:03 Meziani Shift Summary
  • 218155 12/03/07 05:27 Meziani
  • 218150 12/03/07 05:11 Meziani Comparison of Pb runs between BPMs x at -2 and -1
  • 218136 12/03/07 04:04 Meziani HV info for left arm
  • 218135 12/03/07 04:04 Meziani HV info for right arm
  • 218133 12/03/07 04:01 Meziani Compton/OTR Charts
  • 218132 12/03/07 04:01 Liyang Jiang Target Charts
  • 218131 12/03/07 04:00 Meziani BPM/BCM charts
  • 218130 12/03/07 03:59 Meziani Ozone /radiation monitors
  • 218128 12/03/07 03:57 Current Char Current chart
  • 218127 12/03/07 03:56 Meziani Hall A tools
  • 218103 12/03/07 02:37 Liyang Jiang Target Charts
  • 218090 12/03/07 01:28 Xinhu Yan target charts
  • Sunday

  • 218063 12/02/07 23:57 Seonho Choi Swing shift summary
  • 218061 12/02/07 23:52 Seonho Choi L-HRS momentum change
  • 218046 12/02/07 22:41 Seonho Choi HPR momentum change
  • 218027 12/02/07 21:50 yez screen grub 2
  • 218026 12/02/07 21:48 yez screen grub
  • 218025 12/02/07 21:45 Seonho Choi Compton Rates
  • 218024 12/02/07 21:44 Seonho Choi BCM/BPM Strip chart
  • 218022 12/02/07 21:43 Seonho Choi Ozone/Radiation Strip Chart
  • 218020 12/02/07 21:43 Seonho Choi Hall A Tools Page.
  • 218015 12/02/07 21:23 Seonho Choi L-HRS momentum change
  • 218000 12/02/07 20:06 Seonho Choi HRS Momentum Change
  • 217979 12/02/07 19:05 Seonho Choi Target stuck and moving again
  • 217944 12/02/07 16:41 V. Sulkosky Compton DAQ Restarted
  • 217939 12/02/07 16:10 Seonho Choi RF Problem
  • 217938 12/02/07 16:01 Boeglin Shift Summary
  • 217921 12/02/07 15:01 A. Camsonne NaI status
  • 217919 12/02/07 14:53 J.Lopez Target Loop1
  • 217905 12/02/07 13:56 Boeglin BCM/BPM
  • 217903 12/02/07 13:55 Boeglin Ozone/Radiation/Compton
  • 217902 12/02/07 13:54 Boeglin Hall A Tools
  • 217893 12/02/07 13:18 J. Lopez Target snapshot
  • 217818 12/02/07 08:02 Meziani Shift Summary
  • 217800 12/02/07 06:57 Jack Gas system
  • 217795 12/02/07 06:42 R. Michaels N2 replaced, start data taking
  • 217792 12/02/07 06:13 R. Michaels gas update
  • 217791 12/02/07 05:33 R. Michaels gas pressure low ?
  • 217785 12/02/07 05:14 Meziani Targets charts
  • 217782 12/02/07 04:53 Meziani Compton/OTR
  • 217781 12/02/07 04:52 Meziani HVs for left arm
  • 217780 12/02/07 04:51 Meziani HV for Right Arm
  • 217776 12/02/07 04:43 meziani BPM/BCM charts
  • 217775 12/02/07 04:42 Meziani Ozone/Radiation Monitor charts
  • 217767 12/02/07 04:07 R. Michaels start-of-run correction
  • 217726 12/02/07 01:16 R. Michaels error: wrong target
  • 217721 12/02/07 01:08 A. Camsonne NaI HV change
  • 217714 12/02/07 00:34 R. Michaels raster adjustment
  • 217707 12/02/07 00:00 Boeglin Shift Summary
  • Saturday

  • 217671 12/01/07 20:59 yez screen grab for loop3 &cryostat
  • 217668 12/01/07 20:56 yez Screen grab
  • 217667 12/01/07 20:52 Boeglin BCM/BPM/Compton
  • 217666 12/01/07 20:52 Boeglin Ozone/Radiation strip charts
  • 217665 12/01/07 20:51 Adaq Hall A Tools
  • 217624 12/01/07 18:22 yez Correction for run 21362&3161
  • 217613 12/01/07 17:58 yez move target from He4 to Pb
  • 217600 12/01/07 17:33 yez Move target from Pb to He4
  • 217594 12/01/07 17:22 yez move target from H2 to Pb
  • 217593 12/01/07 17:16 yez move target from Dummy to H2
  • 217585 12/01/07 16:39 A. Camsonne Raster current plots
  • 217584 12/01/07 16:39 A. Camsonne Raster current plots
  • 217579 12/01/07 16:25 yez move target from C12 to Fe
  • 217576 12/01/07 16:21 yez Move Dummy to C12 Left
  • 217573 12/01/07 16:10 J. LOPEZ Day Shift Summary
  • 217565 12/01/07 15:53 J. Lopez Ozone snapshot
  • 217556 12/01/07 15:12 J. Lopez Hall A Tools snapshot
  • 217549 12/01/07 14:56 J.Lopez Changing Momentum
  • 217544 12/01/07 14:25 J. Lopez Compton rates snapshot
  • 217542 12/01/07 14:17 Y. Zhang target chart
  • 217536 12/01/07 13:54 J. Lopez BCM and BPM snapshot
  • 217533 12/01/07 13:41 Y. Zhang lead target raster replay
  • 217469 12/01/07 09:30 J. LOPEZ BPM POSITION FOR LAST RUNS
  • 217458 12/01/07 08:09 Z.Meziani Shift summary
  • 217440 12/01/07 07:18 meziani low voltages on S2 PMT 1 and 2
  • 217433 12/01/07 06:52 R. Michaels junk runs, R-arm
  • 217424 12/01/07 06:42 meziani HV left arm
  • 217422 12/01/07 06:40 meziani High voltages Right arm
  • 217421 12/01/07 06:39 meziani BPM/BCM charts
  • 217409 12/01/07 05:41 R. Michaels some wrong start-of-run entries
  • 217398 12/01/07 04:52 meziani Target charts
  • 217389 12/01/07 04:08 meziani Beam raster size run # 3100
  • 217386 12/01/07 04:05 meziani Ozone and Radiation Charts
  • 217385 12/01/07 04:02 meziani OTR/Compton Charts
  • 217382 12/01/07 04:00 meziani Hall A tools
  • 217347 12/01/07 01:22 R. Michaels adjust prescales.
  • 217340 12/01/07 01:09 R. Michaels re: the prescale line in run 3086
  • 217337 12/01/07 01:00 R. Michaels prescale string wrong.
  • 217322 12/01/07 00:01 Boeglin Shift summary